THE ULTRA VIOLET SPECTRUM OF TeO_{2}

\underline{DANIEL F HULLAH} and JOHN M BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ, UK.

The laser induced fluorescence spectrum of TeO_{2} has been recorded between 345 nm and 405, using a high temperature nozzle source. The spectrum shows a considerable improvement in signal-to-noise ratio in comparison with that of Dubois^{1} and this has lead to a reassignment of the vibrational structure, including the band origin. To date, 58 vibrational bands (27 cold and 31 hot) have been reassigned to the ^{1}B_{2}−^{1}A_{1} electronic transition of TeO_{2}. The values of the band origin and the vibrational frequencies for the symmetric stretch (ω_{1}) and bending (ω_{2}) modes are given as follows (values are in cm^{−1})

ν_{0}=25 526  ω′_{1}=678  ω′_{2}=220  ω′′_{1}=822  ω′′_{2}=281.

Measurements at high resolution are planned to allow a rotational analysis. Following on from this work, the products of the high temperature reaction Te + H_{2}O are to be investigated.

Reference

1. I. Dubois, Bulletin de la Societé Royale de Liege, 39^{e} année, n^{o} 1-2, 63 (1970)